Find an equation of the line containing points(-7,9) and parallel to the line 4x=3y+5
Answer by ewatrrr(24785)


Hi,
Using general slope-intercept form for an equation of a line
y = mx + b
line 4x = 3y+5 y = (4/3)x -5/3 Slope is (4/3)
Slope of the line parallel to it will be: 4/3
y = (4/3)x + b using Pt(-7,9) to solve for b
9 = (4/3)*-7 + b
27/3 + 28/3 = b
55/3 = b
y = (4/3)x + 55/3